SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Trigger Operation Meeting (03/18/08)

Attendees: Su Dong, Rahmat Rahmat, Debbie Bard, Kim Hojeong.

Meeting Information

 Date      :  March 18th, 2008
 Time      :  2:00pm - 3:30pm PST
 SLAC Room :  Madrone conf room, ROB 
 
 Dial-in   :  1-510-665-5437, pass code 6649#

--- Minutes ---

Operations

  • Sunday swing: CEN:BIP:L1RATE:AT:ZERO
    Either FCT master crate itself has a problem or Trigger display has a problem. When L1rate is exactly at 0, it can't be a real trigger problem.
  • Saturday swing:
    [22:46] Dataflow trimmed level 3 node number 13, name: bfo-farm115 
         Received  reason 13 via node 3
    [22:46] Level 3 process crashed or aborted. Probable node trim 
         Daemon bfo-farm115 reported error state: SIGNALED
    [23:05] lost connection with FCT epics channels. Restarting FCT gateway didn't help
    
    It was a non-trigger-related network problem, fixed by Steffan & Matthias.

EMT

  • Mark is away from this Thursday until April 8th. He'll be reachable via email.

DCT

GLT

Online

L3

DQM

  • To make sure DCH O2 level problem, we need to check OPR monitoring plots. L1 fast mon is not sensitive enough to show the effect. Need to check TSF weight=3, which is the most sensitive. wieght=2 wouldn't show the effect much either. Should pay attention to the absolute scale. Rahmat will check and report next week.

Trickle injection monitoring

Trigger Configuration

AOB


This page was created by Kim Hojeong; however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.

Last Modified: March 18th, 2008.